OpenStudy (anonymous):

solve the equation x-8sqrtx=-15

OpenStudy (anonymous):

\[x-8\sqrt{x=-15}\]

OpenStudy (mertsj):

\[x-8\sqrt{x}=-15\]

OpenStudy (anonymous):

yesssss

OpenStudy (mertsj):

\[-8\sqrt{x}=-x-15\]

OpenStudy (mertsj):

\[(-8\sqrt{x})^{2}=(-x-15)^{2}\]

OpenStudy (mertsj):

\[64x = x ^{2}+30x +225\]

OpenStudy (anonymous):

Let sqrt(x) = b b^2 - 8b = -15 b^2 - 8b + 15 = 0 (b-3) (b-5) = 0 b = 3, 5 x = 9, 25

OpenStudy (mertsj):

\[x ^{2}-34x + 225 = 0\]

OpenStudy (mertsj):

\[(x-9)(x-25)=0\]

OpenStudy (mertsj):

x=9 or x = 25

OpenStudy (anonymous):

\[2x ^{3}-8x/4x-x ^{2}\]

OpenStudy (anonymous):

can you help me on this?

OpenStudy (mertsj):

Factor the numerator and denominator